What is the probability of a family having 2 girls and 1 boy?
Answer #1

The answer will be 3/8. The chance for a boy or girl is 1/2 as we know mother always gives X and father can give either X or Y. The chance for 2 girl and boy will be calculated as below:

Probability of Girl, Girl and Boy = 1/2 X 1/2 X 1/2 = 1/8

This arrangement could have three possibilities: 1. Girl, Girl, boy 2. Girl, Boy, Girl 3. Boy, Girl, Girl. Thus probability for 2 girl and 1 boy = 1/8 X 3 = 3/8.